OpenStudy (anonymous):

FACTORING: Can x^3+4x^2+24 be factored to find zeros? (fundamental theorem of algebra)

OpenStudy (anonymous):

first find the factors of prime factors of 24

OpenStudy (anonymous):

the factors of 24 are 1,2,3,4,6,8,12 put these values with +ve and -ve sign in the given expression and check for which particular values the expression vanishes

OpenStudy (anonymous):

what's ve

OpenStudy (whpalmer4):

positive (+ve) and negative (-ve)

OpenStudy (anonymous):

by +ve i mean positive sign

OpenStudy (anonymous):

Ohh!

OpenStudy (anonymous):

so do I plug it in the x? and ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and check if the value of the expression vanishes

OpenStudy (anonymous):

so if it gives me zero? would that number then be one of my "solutions"?

OpenStudy (whpalmer4):

Yes, any value of x such that p(x) = 0 is a solution/root/zero.

OpenStudy (anonymous):

then the value of x for which the above expession vanishes will be one of the roots of the equation

OpenStudy (anonymous):

I understand now, both of you, thanks for your kind time (:

OpenStudy (anonymous):

thanks but if you ar not able to find any such specific x by hit and trial method than opt for Cardans method for solving cubic equation

OpenStudy (whpalmer4):

well, Cardan's method isn't what's being requested here — problem asks if it can be factored.

OpenStudy (anonymous):

I found no zeros so I guess that means no real solution?

OpenStudy (anonymous):

since its cubic it must have at least one real solution as imaginary roots occur in conjugate pairs

OpenStudy (anonymous):

can you plz check the sign of 24 its +ve or -ve

OpenStudy (anonymous):

negative I suppose

OpenStudy (anonymous):

but if its negative u have mentioned +ve

OpenStudy (rsadhvika):

\(x^3+4x^2\color{red}{-} 24\)

OpenStudy (rsadhvika):

like that ?

OpenStudy (anonymous):

none of them work, negative or positive factors of 24

OpenStudy (anonymous):

if its x^3+4x^2 - 24 then the expression vanishes for 2 if it vanishes for 2 then x-2 is afactor

OpenStudy (anonymous):

the only thing I close to a 0 is -4 input but only with x^3+4x^2

OpenStudy (anonymous):

@icetea plz check if ur question is x^3+4x^2 + 24 or x^3+4x^2 - 24

OpenStudy (anonymous):

I wrote down -24 but I typed +24 woops, yes it's +24

OpenStudy (rsadhvika):

then the answer would be : It cannot be factored.

OpenStudy (anonymous):

@matricked is right there are roots but they are all decimals so it can't be factored

OpenStudy (anonymous):

I hate 314x^3+1256x^2-7536 then I factored it to 314(x^3+4x^2+24)

OpenStudy (anonymous):

*had

OpenStudy (rsadhvika):

hey u flipped the - sign in the end !

OpenStudy (anonymous):

you can't do that

OpenStudy (anonymous):

so 314x^3+1256x^2-7536 then 314(x^3+4x^2-24)

OpenStudy (anonymous):

there are real roots to this equation and it can be factored keep trying!

OpenStudy (anonymous):

=314(x^3+4x^2-24) =314( x^3 -2x^2 +6x^2 -12x +12x -24 ) =314( x^2(x-2) + 6x(x-2) +12(x-2)) =314((x-2)

OpenStudy (anonymous):

OH I see what you did hmm so then +2 would be my one real root?!

OpenStudy (anonymous):

=314(x^3+4x^2-24) =314( x^3 -2x^2 +6x^2 -12x +12x -24 ) =314( x^2(x-2) + 6x(x-2) +12(x-2)) =314(x-2)(x^2+6x+12)

OpenStudy (anonymous):

i hope the rest u will be able to do...

OpenStudy (whpalmer4):

yes, x= 2 is the real root, and you've got a pair of complex roots

OpenStudy (anonymous):

-12 and +2 yess?

OpenStudy (anonymous):

there are three roots so write the other one.

OpenStudy (whpalmer4):

no. \[x^2+6x+12 =0\] \(a=1,b=6,c=12\) \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-6\pm\sqrt{6^2-4(1)(12)}}{2} = \frac{-6\pm\sqrt{36-48}}{2}=\]

OpenStudy (anonymous):

Your -12 is incorrect

OpenStudy (anonymous):

I thought you only used the discriminant?

OpenStudy (whpalmer4):

the discriminant only allows you to analyze the kinds of roots you'll encounter, not find them in the general case...

OpenStudy (anonymous):

No... you see the \[\pm \] that whpalmer4 placed there.. You suppose to evaluate it at + and at -......

OpenStudy (jhannybean):

If 2 is a zero, why do you get x-2 as a ...factor? LCM = x or something? Or do we just apply this rule for long division standards?

OpenStudy (rsadhvika):

factor theorem

OpenStudy (anonymous):

ooh then the -6+/-sqrt -6 is my complex root yes?

OpenStudy (anonymous):

if 2 is factor then the expresion is divisible by x-2

OpenStudy (anonymous):

you solve for x so x = 2 which is a root according to what matricked stated.

OpenStudy (jhannybean):

I see... and so if a problem has multiple roots, they'd be written such as x - #?

OpenStudy (anonymous):

if it can be factored yes

OpenStudy (jhannybean):

Then using long division with these factors, can see if it equals zero at the end?... trying to remember that root theorem.

OpenStudy (rsadhvika):

once you have roots, you can factor it

OpenStudy (whpalmer4):

@icetea \[\frac{-6\pm\sqrt{-12}}{2} = \frac{-6\pm i2\sqrt{3}}{2} = -3\pm i\sqrt{3}\]

OpenStudy (jhannybean):

I understand that.

OpenStudy (anonymous):

P(x)= (x-a)Q(x) +R(x)

OpenStudy (rsadhvika):

that was an interesting question why (x-p) factors out when p is a root

OpenStudy (anonymous):

@whpalmer that slipped my mind! thanks !!!!!

OpenStudy (jhannybean):

Thank you.

OpenStudy (whpalmer4):

(x-p) factors out when p is a root because P(p) = 0 and P(x) can be written as (x-p)(x-q)... by the factor theorem

OpenStudy (anonymous):

welcome

OpenStudy (jhannybean):

I'll have to look up the factor theorem.

OpenStudy (rsadhvika):

have to lay back and think how factor theorem works

OpenStudy (anonymous):

lol yea lets do that

OpenStudy (whpalmer4):

well, the roots are the spots where f(x) = 0, right? if we write the polynomial as \[(x-r_1)(x-r_2)...(x-r_n) = 0 \text{ (at a zero)}\]then we see our remainder is identically 0 at a root and we can divide off a root as a factor

OpenStudy (rsadhvika):

like when polynomial cuts the x-axis at p, then p is a zero. when u factor out the polynomial :- \(ax^n + bx^{n-1} + cx^{n-2} ... = a(x-p)(x-q).... \)

OpenStudy (rsadhvika):

so we need to think reverse, when p is a zero, it factors into (x-p) i get it !!!